the lines 3x-4y =4 and 6x-8y -7=0 are tangets to the same circle then its radius is
(a) 1/4
(b)1/2
(c)3/4
(d)3/2

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
given lines are 3x-4y=4...........(1),6x-8y=7...........(2)
multiplying eqn 1 by 2........6x-8y=8 and 6x-8y=7
Given lines are parallel lines so 2r=distance between them
2r=(8+7)/sqrt(62+82)
2r=15/10
r=3/4
Thanks and regards
Shaik Aasif
askIITians faculty